Last one...for today ugh
The combined land area of the countries A and B is 173,973 square kilometers. Country A is larger by 773 square kilometers. Determine the land area of each country.
Land area A
Land area B
Thanks in advance. I got through 3 equations by the last set of notes. These word problems will get the best of me.

Answer by rothauserc(4718)

The problem states:
The combined land area of the countries A and B is 173,973 square kilometers. Country A is larger by 773 square kilometers. Determine the land area of each country.
Land area A
Land area B
let x = land area of country B, then we have
x+773+x = 173973
2x = 173200
x = 86600 which is land area of country B
land area of country A = 86600 + 773 = 87373